OpenStudy (anonymous):

Find the equation of a line containing the given point and having the given slope. (2,8), m=5/4 y=___x+___ I'm having problems with the fractions themselves.

OpenStudy (anonymous):

Using the line-slope notation, we can say that \(\Large y = mx + c\). We have x, y and m, so can we work out c?

OpenStudy (anonymous):

y=5/4x+11/2

OpenStudy (anonymous):

I was going more for the concept than the answer, 14yamaka.

OpenStudy (anonymous):

c=y-mx

OpenStudy (anonymous):

sorry :S

OpenStudy (anonymous):

\[y-y_1 = m(x-x_1) \] if I remember from A-Levels lol. So we have \[y-8 = \frac{5}{4}(x-2)\] \[y = \frac{5}{4}x +8- \frac{10}{4}\] \[y=\frac{5}{4}x +\frac{11}{2}\]

OpenStudy (anonymous):

Thanks, guys. Those pesky fractions. Looks like I am really going to have to brush up on my basics. :(

OpenStudy (anonymous):

The derivation of the equation I quoted comes from the general fact that \[m = \frac{y_2-y_1}{x_2-x_1}\] which can be proved easily once one knows the definition of gradient (distance up divided by distance along). We let \[(x_2,y_2)\] be an arbitrary point, and in doing this we can find the gradient of ANY line via: \[m = \frac{y-y_1}{x-x_1}\] which we rearrange to get the equation I began my last answer to...